# Fisher information flow in artificial neural networks

### v1

@@ Abstract
The estimation of continuous parameters from measured data plays a central role in many fields of physics. A key tool in understanding and improving such estimation processes is the concept of Fisher information, which quantifies how information about unknown parameters propagates through a physical system and determines the ultimate limits of precision. With Artificial Neural Networks (ANNs) gradually becoming an integral part of many measurement systems, it is essential to understand how they process and transmit parameter-relevant information internally. Here, we present a method to monitor the flow of Fisher information through an ANN performing a parameter estimation task, tracking it from the input to the output layer. We show that optimal estimation performance corresponds to the maximal transmission of Fisher information, and that training beyond this point results in information loss due to overfitting. This provides a model-free stopping criterion for network training-eliminating the need for a separate validation dataset. To demonstrate the practical relevance of our approach, we apply it to a network trained on data from an imaging experiment, highlighting its effectiveness in a realistic physical setting.
@@ I Introduction
In physics, artificial neural networks (ANNs) are used in a wide variety of research areas, ranging from optics [1] to high energy physics [2], the computation of material parameters [3], the design of experiments [4] and the solution of complex inverse problems [5]. In the context of ANN training, one of the central issues is to understand how models process the information that they receive as input [6, 7, 8, 9].
@@
One particularly successful approach is based on the notion of mutual information (MI) and the information bottleneck method [10, 11, 12], which provide a peek into the training process of the model. MI is also used for regularizing the optimization [13, 14] and for deriving generalization bounds [15, 16]. Nevertheless, this line of research is severely hindered by the challenges of estimating MI in high dimensional spaces [17] and by the issue that the MI diverges for continuous random variables when they are connected by a deterministic transformation [18, 19, 20, 21].
@@
The starting point for the current work is the observation that a whole class of estimation problems in physics centers around the concept of Fisher information (FI) [22, 23]. Arising from the field of (statistical) estimation theory [22], the FI is the key quantity when dealing with estimating continuous parameters from noisy data. The amount of FI one has available on a given parameter ultimately determines how precisely the value of this parameter can be estimated and thus bounds the achievable performance of the estimating model. The FI applies by definition to the estimation of continuous deterministic variables and thus naturally avoids the divergence that occurs for the MI in deterministic ANNs. Parameter estimation tasks of this kind are particularly relevant in optics and photonics, such as in imaging [24, 25, 26, 27], particle levitation [28, 29], classical and quantum metrology [30, 31, 32, 33], to name just a few.
@@
With the increased use of ANNs to solve these parameter estimation problems, it is thus of crucial importance to understand how the FI (rather than the MI) behaves as the data is being processed by the network. Insights in this direction will help us to understand the limitations of ANNs in solving estimation tasks; they can also be expected to provide an essential tool for the design of new network architectures, which perform better in terms of the preservation of information during data processing. Consider as a specific example that in many scenarios, where a parameter is estimated from high-dimensional image data, the data dimensionality needs to be strongly reduced. Analyzing the FI may reveal how an ANN solves the non-trivial task of preserving information during this compression process.
@@
Despite the relevance of the FI in ANNs, the analysis of the FI flow through the network’s layers has, to the best of our knowledge, not been provided yet. Previous works [34, 35, 36, 37, 38] have focused on the FI of an ANN’s parameters such as its weights and biases, which is conceptually different from the FI about parameters inherent to the data itself. Moreover, these earlier works invoke assumptions about the structure of the ANNs or the shape of the distributions. In practice, however, modeling the distributions of the data that is fed into an ANN is very difficult, as the network can generate high-dimensional and arbitrarily complicated representations of the data [39, 40]. The dimensionality of the data inside a layer of the network is equal to the number of the layer’s nodes, which is often of the order of several hundreds even in comparably small models. The FI however, is a functional of the probability density function (PDF) describing the data distribution and there is no feasible way of estimating such a high-dimensional density when it does not take on a very simple form (such as a multivariate Gaussian). In other words, evaluating how much FI is contained in each layer of the ANN is a challenging problem.
@@
Figure 1: A beam of light, originating from the source on the left, is directed onto the target object – the image of a “space invader” placed on a slide orthogonal to the beam. The intensity of the light field transmitted through an objective to a camera carries FI about the parameter of interest $\theta$, which is here the horizontal position of the target object. The image data collected in the experiment is fed into an ANN, which is trained to make a prediction $\hat{\theta}$ of the parameter $\theta$ implicitly contained in the data. For the neural network to achieve satisfactory estimates of $\theta$, the FI must be preserved as much as possible while passing through the ANN from input to output layer.
Here, we solve this problem by using the concept of “linear Fisher information” [41, 42] (LFI) for estimating the true FI directly from the data set – on all of the network’s layers and for arbitrarily complicated distributions. To demonstrate the practical relevance of this approach, we study an ANN trained on data from an optical imaging experiment performed under strong noise conditions. At different stages of the training, we observe how the FI behaves as the data is processed by the network. Using the LFI analysis, we are able to demonstrate how the FI flows through an ANN from input to output. When being well trained, the ANN is able to preserve the FI even when considerable dimensionality reduction takes place. We find that the ANN performs near the ultimate limit of precision that is determined by the physical noise within the measured data. Based on this insight, we are able to show that by calculating the FI contained in the ANN’s input data, we can determine – using only the training data and no validation data – when the network starts to overfit in the training process.
@@ II.1 Imaging experiment
To perform the proposed FI analysis with a practical parameter estimation task, we consider an imaging experiment subject to strong noise, see Fig. 1. The data collected during the experiment is used to train an ANN, which we analyze using FI.
The parameter of interest is the horizontal position of an object (“space invader” etched in an opaque slide), that can be freely translated in the plane of the slide. We illuminate it with broadband light and collect a series of images at different positions of the object, which we then use to train the ANN to estimate this position (see Appendix B for details about the experimental setup). We attenuate the incident light up to a level when the camera noise becomes dominant over the image signal, as shown in Fig. 2 (resulting SNR =2/15=2/15), which leads to a sizeable uncertainty of the position estimates. The statistical distribution of the intensity values measured by each camera pixel is non-trivial (see Appendix B.3), preventing us from constructing an efficient estimator of the position using an analytical expression. The ultimate precision at which this position can be estimated is given by the FI of the input data probability distribution. We calculate the FI of the hidden layers of the network and demonstrate that in the process of training, it gains the ability to transfer FI from the input to the output, thereby achieving this limit.
@@ II.2 Fisher information and the Cramér-Rao lower bound
where the average $\langle\ldots\rangle$ is taken with respect to the data statistics. A function $\hat{\theta}=\tau(X)$ that assigns a prediction $\hat{\theta}$ of the parameter $\theta$ to the data is referred to as an estimator [22]. Restricting ourselves here to the class of unbiased estimators (where $\langle\hat{\theta}-\theta\rangle=0$ holds), the FI bounds the achievable precision of the estimate [44, 23] by
@@
with $\sigma^{2}$ being the variance of $\hat{\theta}$. While this Cramér-Rao lower bound (CRLB) can be generalized also to multiple parameters [43] we focus here only on the single parameter case, for the sake of simplicity.
Figure 2: Camera images of a “space invader” illuminated by broadband light. The position of the object under study is adjusted by moving the slide to which it is attached (in horizontal direction). The leftmost position is shown in the top row (a),(b) and the rightmost position in the bottom row (c),(d). The left column (a),(c) shows the camera image under bright illumination and the right column (b),(d) under weak illumination conditions. The digital numbers (DN) that the camera outputs for each column are shown in the grayscale bars at the bottom.
@@ II.3 Fisher information inside deep neural networks
The ANN we use for the parameter estimation task can be seen as a non-linear function $\tau(X)$ combining multiple functions from the layers of the network [39, 40]. We feed a set of data to the network and extract the transformed data at the consecutive layers. In a feed forward neural network architecture without skip connections, the transformed data in the hidden layers form a Markov chain [45, 10, 46]. The so-called Fisher information inequality [45], a special case of the data processing inequality, assures that the FI does not increase in a Markov chain, i.e., it is a non-increasing function of the ANN’s layer depth [45, 23]. The causes for the loss of FI in ANNs are discussed in Appendix A.1.
Evaluating Eq. (1) by inferring the PDF $p(X;\theta)$ is not feasible, as the random variables we encounter in the imaging experiment are high-dimensional and conventional methods of probability density estimation are hampered by the curse of dimensionality, i.e., the complexity of the problem grows rapidly with increasing dimension. To circumvent this problem, we use the linear Fisher information (LFI) [47, 48, 41, 23, 49, 42, 50, 51], a lower bound of the FI, as an auxiliary quantity. We then employ an auxiliary non-linear network to transform the data in a way that its LFI approximately matches the true FI. The LFI comes with the advantage that it can be precisely estimated from much smaller statistical samples as compared to the FI.
@@ II.4 Linear Fisher information
where the expectation value $\langle$$\rangle$ is taken with respect to the likelihood $p(X;\theta)$. The vector $\bm{\mu}$ is the distribution mean and $\mathbf{\Sigma}$ its covariance matrix. Note that for random variables in a single dimension, this inequality reduces to $J(\theta)=(\partial_{\theta}\mu)^{2}/\sigma^{2}$, with $\sigma^{2}$ being the variance of the random variable. For the one dimensional case we see that the LFI is large when a variation of the parameter $\theta$ induces a large variation of the mean $\bm{\mu}$, while a larger variance implies a smaller LFI. The LFI coincides with the FI, e.g., for a Gaussian distribution with respect to the mean with constant covariance [22] or for an exponential distribution with respect to the rate parameter [22, 52]. What makes the LFI particularly useful is its property of being a lower bound [41, 23, 49] for the FI, i.e., $J(\theta)\leq I(\theta)$.
We deal with the problem of inverting the covariance matrix in Eq. (3), which could be singular, by computing a pseudoinverse instead, based on a principal component analysis. The derivative $\partial_{\theta}\bm{\mu}$ is evaluated with a symmetric finite difference scheme
@@
In order to estimate the LFI from a finite sized data set, the moments $\bm{\mu}$ and $\mathbf{\Sigma}$ can be approximated by the sample mean and the sample covariance matrix, respectively. However, the resulting estimate of the LFI is biased [53], which becomes an issue, e.g., if the step size $\Delta\theta$ is very small. Given that the size of the sample is large enough to accurately estimate the covariance matrix, we compensate for the bias picked up due to a small step size by constructing an approximately unbiased estimator of the LFI (based on [53])
@@
While the FI limits the precision that can be achieved with any possible estimator, the LFI yields a bound when only linear estimators are considered. It can be shown that an estimation at the best possible precision determined by the LFI is locally achieved by the so-called best linear unbiased estimator (BLUE) [22]. The term “locally” refers here to the fact that the BLUE is different for different values of $\theta$. Linear regression does not generally achieve the precision bound determined by the inverse LFI but performs worse. For an ANN it is therefore a non-trivial task to reach the LFI when it is trained on an extended parameter domain.
@@ II.5 LFI maximization
In many cases, the LFI can be used as a good approximation of the FI. An example where the LFI coincides with the FI is the estimation of the unknown mean value from Gaussian and Poissonian distributions. This scenario occurs, e.g., in the scattering of electromagnetic waves [54, 29] where the measurements are on a fundamental level subject to Poissonian noise. However, there are situations where the LFI is significantly smaller than the FI. A well-known example occurs when the estimated parameter is the variance of a normal distribution. When the mean and the variance are independent of each other, the mean does not at all depend on this parameter, i.e., $\partial_{\theta}\bm{\mu}=\mathbf{0}$ and thus $J=0$. In order to solve this problem and present an analysis that applies to arbitrary parameter estimation tasks, we introduce a technique that makes use of the LFI’s behavior under transformations.
@@
We introduce an algorithm that uses non-linear transformations of the data such that the LFI of the transformed data is gradually increased and eventually converges to a value closely matching the true FI. While in previous works [41, 48, 49, 55, 49] bounds on the FI that are tighter than the LFI itself were derived, these methods either rely on model assumptions or are computationally very demanding in high dimensions. These restrictions prevent an application of these approaches to the distributions extracted at the hidden layers of an ANN, which are in general complicated and high-dimensional. Here, in contrast, we present a tool that allows us to construct a suitable bound numerically, using only the data at hand.
@@
The basic idea behind our approach is that we apply a transformation to the data that encodes the same information on multiple channels. While this is very redundant and can, of course, not increase the information content, it increases the information available to the restricted class of linear estimators, which are easy to construct. With more redundancy, more information can be extracted with a linear estimator, hence the LFI is increased. We add more and more redundancy, gradually increasing the LFI, until a convergence criterion is met.
@@
Our starting point to implement this approach is the initial random variable $X$ and the FI $I_{X}$ we want to estimate (the subscript of $I_{X}$ refers to the random variable describing the data, while the explicit dependence on the estimated parameter $\theta$ is omitted for simplicity of notation). We apply a random linear transformation $X^{\prime}=\mathbf{A}X$, followed by a non-linear, component-wise function $Y=\phi(X^{\prime})+\bm{\varepsilon}$ (see Appendix E) where $\bm{\varepsilon}$ is a random vector whose components are small compared to the components of $\phi(X^{\prime})$. The additive term $\bm{\varepsilon}$ acts as a regularizer avoiding extremely narrow distributions. We then calculate the LFI of $Y$, which, for the right transformation, provides a very good approximation of the true FI $I_{X}$. When the rank of $\mathbf{A}$ is equal to the dimension $d$ of $X$, $\phi$ is bijective, and moreover, when the information reduction from the added noise is small, there is no channel for a significant loss of FI in the transformation (see Appendix A.1). For dimensions $d^{\prime}$ of $X^{\prime}$ greater than $d$, we observe for the distributions involved that the BLUE is in most cases able to extract more information from the transformed data compared to the original one. As a result, the LFI will likely increase. This behaviour can be seen, e.g., in Fig. 3. Randomly embedding the data into a high-dimensional space and constructing a BLUE from the transformed data is analogous to an extreme learning machine [56, 57] that is trained locally on a small domain of parameters $\theta$. The estimated LFI of the transformed data is thus an estimate of the amount of information that can be extracted with an extreme learning machine. Note that this auxiliary network cannot be used to perform the original estimation task on the whole parameter domain, as it would constitute a suitable estimator only in a small domain around a given parameter value. This is also one of the reasons why our method is computationally feasible, since constructing such a local estimator is a much simpler task than constructing a single estimator that performs well on an extended parameter domain.
@@
Merely applying a single random embedding into a higher dimensional space is typically not a sufficient strategy. This is because the dimension of the transformed random variable can be too small (resulting in an underestimation of the FI) or too large (resulting in an imprecise estimate of the LFI due to the finite size of the data set). Employing a series of embeddings with a gradually increasing dimension allows us to calculate the LFI of $X^{\prime}$ as a function of $d^{\prime}$ and analyze the convergence of the process. We expect a monotonic curve with a plateau, i.e., a dimension, after which the increase of $J_{X^{\prime}}$ is small compared to its increase in a region near $d^{\prime}=d$. Since our estimator of the LFI is unbiased when sufficient data is used to estimate the LFI precisely (see Eq. 5), our LFI curve is an unbiased estimate of the ’true’ curve, which cannot grow without bounds. Moreover, using the methods in Appendix A.2 we can access the statistical properties of the estimated LFI curve as a function of the dimension, the finite-difference step size and the sample size. Specifically, Eq. 8 allows us to estimate the error we make by approximating the true LFI curve with our LFI curve using a finite sample. We speak of reaching the plateau when our estimated LFI curve enters a regime of very slow growth as a function of $d^{\prime}$, given that our estimated error remains small as $d^{\prime}$ is varied. If the estimated error as a function of $d^{\prime}$ grows too large before the plateau of the LFI curve is observed, the error can be reduced by increasing the step size or the sample size. If this is not possible, e.g., due to limitations of the experimental setup, better convergence can be achieved by measuring more data and increasing the sample size.
@@
One scenario, where we use an alternative approach, is when the LFI is already very close to the FI. In such a case we do not observe an increase of the LFI for increasing $d^{\prime}$. We then use the LFI averaged over the dimensions as the estimate for the FI. We emphasize here that, as an alternative to the extreme learning machine as the auxiliary network, one can also use other types of neural networks, with the objective of maximizing the LFI, e.g., with gradient descent. However, the loss functions that are minimized in order to identify an optimal set of weights are then typically non-convex. This may result in the algorithm becoming trapped in a local minimum.
@@
Our algorithm converges to the true FI, when the BLUE of the transformed data extracts all of the FI that is available in the original data, when the parameter is known to be in a small region around a parameter of reference. Satisfying this condition is equivalent to finding an estimator that locally achieves the (biased version of the) CRLB, around this reference parameter. As shown in [58], such an estimator exists, even for non-exponential families of distributions. Since extreme learning machines are universal function approximators [57], for large dimensions of the transformed data $d^{\prime}$, the auxiliary network is in principle able to fit this estimator and thus to perform at this limit. Assuming that the size of the data set allows the LFI of the transformed data to be accurately estimated, it matches the true FI to a good approximation if $d^{\prime}$ is large enough so that the corresponding extreme learning machine has a sufficient complexity to fit, e.g., the estimator presented in [58].

### v2

@@ Abstract
The estimation of continuous parameters from measured data plays a central role in many fields of physics. A key tool in understanding and improving such estimation processes is the concept of Fisher information, which quantifies how information about unknown parameters propagates through a physical system and determines the ultimate limits of precision. With artificial neural networks (ANNs) gradually becoming an integral part of many measurement systems, it is essential to understand how they process and transmit parameter-relevant information internally. Here, we present a method to monitor the flow of Fisher information through an artificial neural network performing a parameter estimation task, tracking it from the input to the output layer. We show that optimal estimation performance corresponds to the maximal transmission of Fisher information, and that training beyond this point results in information loss due to overfitting. This provides a model-free stopping criterion for network training-eliminating the need for a separate validation dataset. To demonstrate the practical relevance of our approach, we apply it to a network trained on data from an imaging experiment, highlighting its effectiveness in a realistic physical setting.
@@ I Introduction
In physics, artificial neural networks (ANNs) are used in a wide variety of research areas, ranging from optics [1] to high-energy physics [2], the computation of material parameters [3], the design of experiments [4], and the solution of complex inverse problems [5]. In the context of ANN training, one of the central issues is to understand how models process the information that they receive as input [6, 7, 8, 9].
@@
One particularly successful approach is based on the notion of mutual information (MI) and the information bottleneck method [10, 11, 12], which provide a peek into the training process of the model. MI is also used for regularizing the optimization [13, 14] and for deriving generalization bounds [15, 16]. Nevertheless, this line of research is severely hindered by the challenges of estimating MI in high-dimensional spaces [17] and by the issue that the MI diverges for continuous random variables when they are connected by a deterministic transformation [18, 19, 20, 21].
@@
The starting point for the current work is the observation that a whole class of estimation problems in physics centers around the concept of Fisher information (FI) [22, 23]. Arising from the field of (statistical) estimation theory [22], FI is the key quantity when dealing with estimating continuous parameters from noisy data. The amount of FI one has available on a given parameter ultimately determines how precisely the value of this parameter can be estimated and, thus, bounds the achievable performance of the estimating model. As FI applies by definition to the estimation of continuous deterministic variables it naturally avoids the divergence that occurs for the MI in deterministic ANNs. Parameter estimation tasks of this kind are particularly relevant in optics and photonics, such as in imaging [24, 25, 26, 27], particle levitation [28, 29], and classical and quantum metrology [30, 31, 32, 33], to name just a few.
@@
With the increased use of ANNs to solve these parameter estimation problems, it is, thus, of crucial importance to understand how FI (rather than MI) behaves as the data are being processed by the network. Insights in this direction will help us to understand the limitations of ANNs in solving estimation tasks; they can also be expected to provide an essential tool for the design of new network architectures, which perform better in terms of the preservation of information during data processing. Consider as a specific example that in many scenarios, where a parameter is estimated from high-dimensional image data, the data dimensionality needs to be strongly reduced. Analyzing the FI may reveal how an ANN solves the nontrivial task of preserving information during this compression process.
@@
Despite the relevance of FI in ANNs, the analysis of the FI flow through the network’s layers has, to the best of our knowledge, not been provided yet. Previous works [34, 35, 36, 37, 38] have focused on the FI of an ANN’s parameters such as its weights and biases, which is conceptually different from the FI about parameters inherent to the data itself. Moreover, these earlier works invoke assumptions about the structure of the ANNs or the shape of the distributions. In practice, however, modeling the distributions of the data that are fed into an ANN is very difficult, as the network can generate high-dimensional and arbitrarily complicated representations of the data [39, 40]. The dimensionality of the data inside a layer of the network is equal to the number of the layer’s nodes, which is often of the order of several hundreds even in comparably small models. The FI, however, is a functional of the probability density function describing the data distribution, and there is no feasible way of estimating such a high-dimensional density when it does not take on a very simple form (such as a multivariate Gaussian). In other words, evaluating how much FI is contained in each layer of the ANN is a challenging problem.
@@
Figure 1: A beam of light, originating from the source on the left, is directed onto the target object – the image of a “space invader” placed on a slide orthogonal to the beam. The intensity of the light field transmitted through an objective to a camera carries FI about the parameter of interest $\theta$, which is here the horizontal position of the target object. The image data collected in the experiment are fed into an ANN, which is trained to make a prediction $\hat{\theta}$ of the parameter $\theta$ implicitly contained in the data. For the neural network to achieve satisfactory estimates of $\theta$, the FI must be preserved as much as possible while passing through the ANN from input to output layer.
Here, we solve this problem by using the concept of “linear Fisher information” [41, 42] (LFI) for estimating the true FI directly from the dataset – on all of the network’s layers and for arbitrarily complicated distributions. To demonstrate the practical relevance of this approach, we study an ANN trained on data from an optical imaging experiment performed under strong noise conditions. At different stages of the training, we observe how the FI behaves as the data are processed by the network. Using the LFI analysis, we are able to demonstrate how FI flows through an ANN from input to output. When being well trained, the ANN is able to preserve FI even when considerable dimensionality reduction takes place. At this point the ANN performs near the ultimate limit of precision that is determined by the physical noise within the measured data. Based on this insight, we are able to show that by calculating the FI contained in the ANN’s input data, we can determine – using only the training data and no validation data – when the network starts to overfit in the training process.
@@ II.1 Imaging experiment
To perform the proposed FI analysis with a practical parameter estimation task, we consider an imaging experiment subject to strong noise, see Fig. 1. The data collected during the experiment are used to train an ANN, which we analyze using FI.
The parameter of interest is the horizontal position of an object (“space invader” etched in an opaque slide), that can be freely translated in the plane of the slide. We illuminate it with broadband light and collect a series of images at different positions of the object, which we then use to train the ANN to estimate this position (see Appendix B for details about the experimental setup). We attenuate the incident light up to a level when the camera noise becomes dominant over the image signal, as shown in Fig. 2 (resulting SNR of around $13\%$), which leads to a sizable uncertainty of the position estimates. The statistical distribution of the intensity values measured by each camera pixel is nontrivial (see Appendix B.3), preventing us from constructing an efficient estimator of the position using an analytical expression. The ultimate precision at which this position can be estimated is given by the FI of the input data probability distribution. We calculate the FI of the hidden layers of the network and demonstrate that in the process of training, it gains the ability to transfer FI from the input to the output, thereby achieving this limit.
@@ II.2 Fisher information and the Cramér-Rao lower bound
where the average $\langle\cdots\rangle$ is taken with respect to the data statistics. A function $\hat{\theta}=\tau(X)$ that assigns a prediction $\hat{\theta}$ of the parameter $\theta$ to the data is referred to as an estimator [22]. Restricting ourselves here to the class of unbiased estimators (where $\langle\hat{\theta}-\theta\rangle=0$ holds), the FI bounds the achievable precision of the estimate [44, 23] by
@@
with $\sigma^{2}$ being the variance of $\hat{\theta}$. While this Cramér-Rao lower bound (CRLB) can be generalized also to multiple parameters [43], we focus here only on the single parameter case, for the sake of simplicity.
Figure 2: Camera images of a “space invader” illuminated by broadband light. The position of the object under study is adjusted by moving the slide to which it is attached (in the horizontal direction). The leftmost position is shown in the top row (a),(b) and the rightmost position in the bottom row (c),(d). The left column (a),(c) shows the camera image under bright illumination and the right column (b),(d) under weak illumination conditions. The digital numbers (DN) that the camera outputs as photons are impinging are shown in the grayscale bars at the bottom.
@@ II.3 Fisher information inside deep neural networks
The ANN we use for the parameter estimation task can be seen as a nonlinear function $\tau(X)$ combining multiple functions from the layers of the network [39, 40]. We feed a set of data to the network and extract the transformed data at the consecutive layers. In a feedforward neural network architecture without skip connections, the transformed data in the hidden layers form a Markov chain [45, 10, 46]. The so-called Fisher information inequality [45], a special case of the data processing inequality, assures that the FI does not increase in a Markov chain, i.e., it is a nonincreasing function of the ANN’s layer depth [45, 23]. The causes for the loss of FI in ANNs are discussed in Appendix A.1.
Evaluating Eq. (1) by inferring the probability density function $p(X;\theta)$ is not feasible, as the random variables we encounter in the imaging experiment are high dimensional and conventional methods of probability density estimation are hampered by the curse of dimensionality; i.e., the complexity of the problem grows rapidly with increasing dimension. To circumvent this problem, we use the LFI [47, 48, 41, 23, 49, 42, 50, 51], a lower bound of the FI, as an auxiliary quantity. We then employ an auxiliary nonlinear network to transform the data in a way that its LFI approximately matches the true FI. The LFI comes with the advantage that it can be precisely estimated from much smaller statistical samples as compared to the FI.
@@ II.4 Linear Fisher information
where the expectation value $\langle\cdots\rangle$ is taken with respect to the likelihood $p(X;\theta)$. The vector $\bm{\mu}$ is the distribution mean and $\mathbf{\Sigma}$ its covariance matrix. Note that, for random variables in a single dimension, this inequality reduces to $J(\theta)=(\partial_{\theta}\mu)^{2}/\sigma^{2}$, with $\sigma^{2}$ being the variance of the random variable. For the one-dimensional case, we see that the LFI is large when a variation of the parameter $\theta$ induces a large variation of the mean $\bm{\mu}$, while a larger variance implies a smaller LFI. The LFI coincides with the FI, e.g., for a Gaussian distribution with respect to the mean with constant covariance [22] or for an exponential distribution with respect to the rate parameter [22, 52]. What makes the LFI particularly useful is its property of being a lower bound [41, 23, 49] for the FI, i.e., $J(\theta)\leq I(\theta)$.
We deal with the problem of inverting the covariance matrix in Eq. (3), which could be singular, by computing a pseudoinverse instead, based on a principal component analysis. The derivative $\partial_{\theta}\bm{\mu}$ is evaluated with a symmetric finite-difference scheme
@@
In order to estimate the LFI from a finite-sized dataset, the moments $\bm{\mu}$ and $\mathbf{\Sigma}$ can be approximated by the sample mean and the sample covariance matrix, respectively. However, the resulting estimate of the LFI is biased [53], which becomes an issue, e.g., if the step size $\Delta\theta$ is very small. Given that the size of the sample is large enough to accurately estimate the covariance matrix, we compensate for the bias picked up due to a small step size by constructing an approximately unbiased estimator of the LFI (based on Ref. [53])
@@
While the FI limits the precision that can be achieved with any possible estimator, the LFI yields a bound when only linear estimators are considered. It can be shown that an estimation at the best possible precision determined by the LFI is locally achieved by the so-called best linear unbiased estimator (BLUE) [22]. The term “locally” refers here to the fact that the BLUE is different for different values of $\theta$. Linear regression does not generally achieve the precision bound determined by the inverse LFI but performs worse. For an ANN, it is, therefore, a nontrivial task to reach the LFI when it is trained on an extended parameter domain.
@@ II.5 LFI maximization
In many cases, the LFI can be used as a good approximation of the FI. An example where the LFI coincides with the FI is the estimation of the unknown mean value from Gaussian and Poissonian distributions. This scenario occurs, e.g., in the scattering of electromagnetic waves [54, 29], where the measurements are on a fundamental level subject to Poissonian noise. However, there are situations where the LFI is significantly smaller than the FI. A well-known example occurs when the estimated parameter is the variance of a normal distribution. When the mean and the variance are independent of each other, the mean does not at all depend on this parameter, i.e., $\partial_{\theta}\bm{\mu}=\mathbf{0}$ and, thus, $J=0$. In order to solve this problem and present an analysis that applies to arbitrary parameter estimation tasks, we introduce a technique that makes use of the LFI’s behavior under transformations.
@@
We introduce an algorithm that uses nonlinear transformations of the data such that the LFI of the transformed data is gradually increased and eventually converges to a value closely matching the true FI. While in previous works [41, 48, 49, 55, 49] bounds on the FI that are tighter than the LFI itself were derived, these methods either rely on model assumptions or are computationally very demanding in high dimensions. These restrictions prevent an application of these approaches to the distributions extracted at the hidden layers of an ANN, which are, in general, complicated and high dimensional. Here, in contrast, we present a tool that allows us to construct a suitable bound numerically, using only the data at hand.
@@
The basic idea behind our approach is that we apply a transformation to the data that encodes the same information on multiple channels. While this is very redundant and can, of course, not increase the information content, it increases the information available to the restricted class of linear estimators, which are easy to construct. With more redundancy, more information can be extracted with a linear estimator; hence, the LFI is increased. We add more and more redundancy, gradually increasing the LFI, until a convergence criterion is met.
@@
Our starting point to implement this approach is the initial random variable $X$ and the FI $I_{X}$ we want to estimate (the subscript of $I_{X}$ refers to the random variable describing the data, while the explicit dependence on the estimated parameter $\theta$ is omitted for simplicity of notation). We apply a random linear transformation $X^{\prime}=\mathbf{A}X$, followed by a nonlinear, componentwise function $Y=\phi(X^{\prime})+\bm{\varepsilon}$ (see Appendix E) where $\bm{\varepsilon}$ is a random vector whose components are small compared to the components of $\phi(X^{\prime})$. The additive term $\bm{\varepsilon}$ acts as a regularizer avoiding extremely narrow distributions. We then calculate the LFI of $Y$, which, for the right transformation, provides a very good approximation of the true FI $I_{X}$. When the rank of $\mathbf{A}$ is equal to the dimension $d$ of $X$, $\phi$ is bijective, and moreover, when the information reduction from the added noise is small, there is no channel for a significant loss of FI in the transformation (see Appendix A.1). For dimensions $d^{\prime}$ of $X^{\prime}$ greater than $d$, we observe for the distributions involved that the BLUE is in most cases able to extract more information from the transformed data compared to the original one. As a result, the LFI will likely increase. This behavior can be seen, e.g., in Fig. 3. Randomly embedding the data into a high-dimensional space and constructing a BLUE from the transformed data is analogous to an extreme learning machine [56, 57] that is trained locally on a small domain of parameters $\theta$. The estimated LFI of the transformed data is, thus, an estimate of the amount of information that can be extracted with an extreme learning machine. Note that this auxiliary network cannot be used to perform the original estimation task on the whole parameter domain, as it would constitute a suitable estimator only in a small domain around a given parameter value. This is also one of the reasons why our method is computationally feasible, since constructing such a local estimator is a much simpler task than constructing a single estimator that performs well on an extended parameter domain.
@@
Merely applying a single random embedding into a higher-dimensional space is typically not a sufficient strategy. This is because the dimension of the transformed random variable can be too small (resulting in an underestimation of the FI) or too large (resulting in an imprecise estimate of the LFI due to the finite size of the dataset). Employing a series of embeddings with a gradually increasing dimension allows us to calculate the LFI of $X^{\prime}$ as a function of $d^{\prime}$ and analyze the convergence of the process. We expect a monotonic curve with a plateau, i.e., a dimension, after which the increase of $J_{X^{\prime}}$ is small compared to its increase in a region near $d^{\prime}=d$. Since our estimator of the LFI is unbiased when sufficient data are used to estimate the LFI precisely [see Eq. (5)], our LFI curve is an unbiased estimate of the “true” curve, which cannot grow without bounds. Moreover, using the methods in Appendix A.2, we can access the statistical properties of the estimated LFI curve as a function of the dimension, the finite-difference step size, and the sample size. Specifically, Eq. (8) allows us to estimate the error we make by approximating the true LFI curve with our LFI curve using a finite sample. We speak of reaching the plateau when our estimated LFI curve enters a regime of very slow growth as a function of $d^{\prime}$, given that our estimated error remains small as $d^{\prime}$ is varied. If the estimated error as a function of $d^{\prime}$ grows too large before the plateau of the LFI curve is observed, the error can be reduced by increasing the step size or the sample size. If this is not possible, e.g., due to limitations of the experimental setup, better convergence can be achieved by measuring more data and increasing the sample size.
@@
One scenario, where we use an alternative approach, is when the LFI is already very close to the FI. In such a case we do not observe an increase of the LFI for increasing $d^{\prime}$. We then use the LFI averaged over the dimensions as the estimate for the FI. We emphasize here that, as an alternative to the extreme learning machine as the auxiliary network, one can also use other types of neural networks, with the objective of maximizing the LFI, e.g., with gradient descent. However, the loss functions that are minimized in order to identify an optimal set of weights are then typically nonconvex. This may result in the algorithm becoming trapped in a local minimum.
@@
Our algorithm converges to the true FI, when the BLUE of the transformed data extracts all of the FI that is available in the original data, when the parameter is known to be in a small region around a parameter of reference. Satisfying this condition is equivalent to finding an estimator that locally achieves the CRLB, around this reference parameter. As shown in Ref. [58], such an estimator exists, even for nonexponential families of distributions. Since extreme learning machines are universal function approximators [57], for large dimensions of the transformed data $d^{\prime}$, the auxiliary network is, in principle, able to fit this estimator and, thus, to perform at this limit. Assuming that the size of the dataset allows the LFI of the transformed data to be accurately estimated, it matches the true FI to a good approximation if $d^{\prime}$ is large enough so that the corresponding extreme learning machine has a sufficient complexity to fit, e.g., the estimator presented in Ref. [58].